# FLASC: A Flare-Sensitive Clustering Algorithm

### v1

@@ Abstract.
We present FLASC, an algorithm for flare-sensitive clustering. Our algorithm builds upon HDBSCAN*—which provides high-quality density-based clustering performance—through a post-processing step that differentiates branches within the detected clusters’ manifold, adding a type of pattern that can be discovered. Two variants of the algorithm are presented, which trade computational cost for noise robustness. We show that both variants scale similarly to HDBSCAN* in terms of computational cost and provide stable outputs using synthetic data sets, resulting in an efficient flare-sensitive clustering algorithm. In addition, we demonstrate the algorithm’s benefit in data exploration over HDBSCAN* clustering on two real-world data sets.
flare detection, density-based clustering, multi-dimensional persistence
copyright: ccccs: Computing methodologies Cluster analysisccs: Computing methodologies Dimensionality reduction and manifold learningccs: Information systems Clustering
Figure 1. Weighted cluster segmentation for HDBSCAN* (left) and FLASC (right). FLASC distinguishes flares in clusters with distinct labels. Point cloud adapted from HDBSCAN*’s online tutorial pages (McInnes et al., 2017, 2022).
Two scatterplots with points coloured by their cluster membership strength. HDBSCAN* detects the Y-shaped cluster as a single subgroup. FLASC detects four subgroups within that cluster: one for each branch and one for the centre where the branches merge.
Figure 1. Weighted cluster segmentation for HDBSCAN* (left) and FLASC (right). FLASC distinguishes flares in clusters with distinct labels. Point cloud adapted from HDBSCAN*’s online tutorial pages (McInnes et al., 2017, 2022).
1. Introduction
Exploratory Data Analysis (EDA)—i.e., searching for interesting patterns in data—is ubiquitous in data science and knowledge discovery workflows. Detecting which subpopulations exist in a data set is a common step in EDA. Not only clusters but also their shapes can reveal relevant subgroups. Specifically, flares—i.e. the branches in a cluster’s manifold—can represent meaningful subpopulations (f.i., see (Lum et al., 2013; Skaf and Laubenbacher, 2022; Kamruzzaman et al., 2018; Reaven and Miller, 1979)). Informally, flares are attachments to a cluster’s core, such as the Y-shaped cluster’s branches in Figure 1.
Several data exploration techniques exist that can detect such branch-based subgroups. Structure learning algorithms—such as Bonnaire et al.’s mixture model approach (Bonnaire et al., 2022) and Mao et al.’ Reversed Graph Embedding (Mao et al., 2017)—model a data set’s principal curve as a graph representing the paths through the middle of the point cloud. These algorithms can, for example, extract road networks from GPS car position samples (Bonnaire et al., 2022) or describe developmental trajectories in single-cell gene expression data (Qiu et al., 2017; Mao et al., 2017). The resulting graphs can also partition the data by the segments between their intersections to detect flare-sensitive subgroups (Chervov et al., 2020). Mapper (Singh et al., 2007) summarises a data set’s structure over manually-specified lens dimensions allowing researchers to inject their domain knowledge into the analysis. The algorithm has been used successfully for finding patterns that previously went unnoticed (Lum et al., 2013; Skaf and Laubenbacher, 2022). When a centrality metric is used as lens dimension, Mapper’s output graph describes the data’s branching structure from which flare-sensitive subgroups can be extracted (f.i. see (Kamruzzaman et al., 2018; Madhobi et al., 2019)).
Clustering algorithms, on the other hand, generally cannot detect this type of subgroup because there is no gap that separates flares from their cluster (Carlsson, 2014). For example, HDBSCAN* (Campello et al., 2013, 2015)—which provides state-of-the-art density-based clustering performance—detects the Y-shaped orange cluster in Figure 1 as a single entity. In the present paper, we extend HDBSCAN* with a flare-detection post-processing step to enable it to detect branch-based subgroups (for example, see Figure 1). This approach provides several benefits: (1) operating on the clusters found by HDBSCAN* reduces the effects of noisy observations, and (2) not building an explicit model of the data’s structure avoids computational costs. We call the resulting algorithm Flare-Sensitive Clustering (FLASC). We empirically analyse its computational cost and stability on synthetic data sets to show that the cost of detecting flares is relatively low. In addition, we demonstrate FLASC on two real-world data sets, illustrating its benefits for data exploration.
In the remainder of this section, we discuss related work and our contributions. Section 2 briefly recaps the HDBSCAN* clustering algorithm. Section 3 describes how FLASC builds on HDBSCAN* to detect branches within clusters and discusses the algorithm’s complexity and stability. Section 4 presents our empirical analyses that show how the algorithm behaves. Finally, Sections 5 and 6 discuss our results and present our conclusions.
@@ 1. Introduction
1.1. Clustering for Exploratory Data Analysis
Currently, HDBSCAN* (Campello et al., 2013, 2015) provides state-of-the-art clustering performance. The algorithm is well suited for exploring unfamiliar data because—unlike older popular clustering algorithms—HDBSCAN* does not require the number of clusters or the distance between clusters to be specified in advance. Additionally, the algorithm is robust to noise and makes no assumptions about the data’s underlying distribution.
HDBSCAN* is a density-based clustering algorithm. Informally, density-based clustering specifies clusters as regions of high density separated by regions of lower density. Visually, this notion is very intuitive. For example, even without the colours, one could easily differentiate the clusters from the noise in Figure 1. In addition, density-based clusters are not limited to convex shapes, and the formulation provides a natural way to separate noise points from clusters.
The principles of density-based clustering were pioneered by Wishart in (1969) (Wishart, 1969). In that article, Wishart presents One Level Mode Analysis, the—to our knowledge—first clustering algorithm based on non-parametric density estimation (as mentioned in (Campello et al., 2015)). Hartigan later formalised the concepts of density contour clusters and density contour trees (Hartigan, 1975), which underpin all density-based clustering algorithms in some form (Campello et al., 2015). The latter concept can also be viewed from a topological perspective, as the split tree constructed in a filtration over the density. See, for example, Figure 2, which shows a 1D continuous density profile with the resulting density contour clusters and density contour tree.
Since HDBSCAN*’s introduction, several studies have implemented and adapted the algorithm: McInnes and Healy (McInnes and Healy, 2017) improved the algorithm’s computational performance by using space trees for finding the data points’ nearest neighbours and provide a popular and efficient Python implementation (McInnes et al., 2017). Stewart and Al-Khassaweneh (Stewart and Al-Khassaweneh, 2022) created a Java implementation with a novel prediction technique for unseen data points. Jackson et al. (Jackson et al., 2018) presented an approximate HDBSCAN* algorithm that uses NN-descent (Dong et al., 2011) for finding the nearest neighbours providing fast distributed performance. Malzer and Baum (Malzer and Baum, 2020) introduced a cluster selection distance threshold that effectively creates a hybrid between DBSCAN’s (Ester et al., 1996) and HDBSCAN*’s cluster selection, improving the algorithm’s performance on data sets with small clusters and a large density variability. Neto et al. (Neto et al., 2021) showed how Relative Neighbourhood Graphs (RNGs) (Toussaint, 1980) can be used to efficiently compute HDBSCAN* cluster hierarchies for multiple min cluster size values. Their follow-up work presented MustaCHE, a visualisation tool for the resulting meta-cluster hierarchy (Neto et al., 2018). To our knowledge, no previous study has adapted HDBSCAN* for detecting flares.
1.2. Flare Detection
@@
As mentioned before, clustering algorithms generally cannot detect flares within clusters. Using Figure 1 as an example, the Y-shaped cluster is found as a single entity by HDBSCAN*. This is expected because no lower-density region separates the flares. In other words, there is a path between data points in different branches that exclusively goes through data points ‘that lie close together’. From a topological perspective, as explained by Carlsson (Carlsson, 2014), clustering algorithms capture the 0-dimensional homology of the data space. That is, clustering describes the connected components in a simplicial complex of the data. Flares are connected in the simplicial complex. Therefore, they have a vanishing homology and cannot be detected as clusters.
It should be noted, though, that implicit in this argument is the assumption that the density across flares is constant (or decreasing). Subpopulations in real-world data sets tend to have some location in feature space where observations are more likely. These locations are detectable as local density maxima forming distinct clusters in a density contour tree, allowing the data points surrounding them to be classified as a particular subgroup. The corresponding cluster hierarchy, however, does not describe the data’s branching structure; it only detects the presence of these local maxima.
Having observed that data sets tend to have a central core connecting everything, Carlsson (Carlsson, 2014) proposes functional persistence to detect branches. The underlying idea is to remove the ‘core’ of the data set, separating branches from each other and making them detectable as clusters (see Figure 3). Carlsson (Carlsson, 2014) provides a centrality measure based on each data point’s distances sum to specify which data points are part of the ‘core’. Here, central data points have lower distances than points towards the extreme ends of the feature space. How much of the ‘core’ should be removed is controlled manually.
@@
Ideally, the complete branching hierarchy is evaluated, describing how branches grow and merge as more of the ‘core’ gets introduced. Note the similarity to HDBSCAN* building the entire cluster hierarchy. Detecting a data set’s clusters and branches in a single filtration requires analysing the multi-parameter persistence over the density and eccentricity, which is expensive to compute (Lesnick and Wright, 2022; Kerber and Rolle, 2021). The resulting bigraded hierarchy is also complicated to work with, as there is no compact representation (Carlsson, 2014) (though research into usable representations is ongoing (Botnan et al., 2022)), and existing visualisations for two-parameter filtrations are non-trivial (Lesnick and Wright, 2015; Scoccola and Rolle, 2023). Strategies that simultaneously vary both dimensions in a single-parameter filtration also exist (Chazal et al., 2009); however, they remain computationally expensive (Vandaele et al., 2021).
@@
It is possible, however, to efficiently compute the complete branching hierarchy at a fixed density threshold using a graph approximation of the data. This replaces the question of how much ‘core’ to remove with which data points should be connected (which we will answer based on HDBSCAN*’s design). Conceptually, this approach can be thought of as creating a sequence of subgraphs which progressively include more and more central points and tracking the remaining connected components. Interestingly, a method like this has been used by Li et al. to detect actual branches in 3D models of plants (Li et al., 2017). Vandaele et al. show that flare-detection in graph approximations works well with a centrality metric based on the maximum shortest path for each point in the network when the graphs accurately approximate the data (Vandaele et al., 2021).
1.3. Contributions
Our main contribution is FLASC: an algorithm that combines density-based clustering and the graph-based flare detection approach described in Section 1.2 to efficiently detect clusters and their flares in unfamiliar data using intuitive parameters and interpretable results. We propose two methods for constructing the approximation graphs with which flares are detected in clusters that naturally arise from HDBSCAN*’s design. We also provide a practical centrality metric for points within a cluster that is computable in linear complexity. Combining density-based clustering and the flare detection approach into a single algorithm provides several attractive properties: (1) the resulting algorithm can deal with data sets that contain more than one cluster, (2) the flare detection sensitivity is increased by operating only on density-based clusters, which suppresses spurious noisy connectivity between branches, and (3) the branching structures are described at each cluster’s density, allowing the algorithm to operate at multiple distance scales.
2. HDBSCAN* Revisited
@@
In this section, we briefly introduce HDBSCAN*, following Campello et al.’s (Campello et al., 2015) explanation. We refer the reader to McInnes and Healy (McInnes and Healy, 2017) for a more formal, statistically motivated description of the algorithm.
Let $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ be a data set consisting of $N$ feature vectors $\mathbf{x}_{(\cdot)}$ and a distance metric $d(\mathbf{x}_{i},\mathbf{x}_{j})$. Then, we define a point’s core distance $\kappa(\mathbf{x}_{i})$ to be the distance to its $k$-th nearest neighbour and the mutual reachability distance to be:
@@
(1) $d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{j})=\begin{cases}\max{\left\{\kappa(\mathbf{x}_{i}),\kappa(\mathbf{x}_{j}),d(\mathbf{x}_{i},\mathbf{x}_{j})\right\}}&\text{if }\mathbf{x}_{i}\neq\mathbf{x}_{j}\text{,}\\ 0&\text{otherwise,}\end{cases}$
where the value of $k$ is specified manually.
@@
Standard Single Linkage Clustering (Sibson, 1973) is applied on the data set with the mutual reachability distance $(\mathbf{X},d_{mreach})$. The resulting clustering hierarchy is then simplified using a manually specified minimum cluster size $m_{c}$. From the root down, only the sides of a split containing more than $m_{c}$ points are considered to represent clusters. Sides with fewer points are interpreted as ‘falling out of the cluster’ or the cluster disappearing completely. The result is a condensed cluster hierarchy that still contains detailed data point membership information.
This condensed hierarchy is expressed in terms of density rather than distance, which is defined as $\lambda_{k}(\cdot)=1/d_{mreach}(\cdot)$. Here, $k$—as used in $\kappa(\cdot)$—acts as a smoothing factor for the density estimation. The relative stability of the clusters is also defined in terms of density. Specifically, cluster $C_{j}$’s relative stability $\sigma_{k}(C_{j})$ is:
(2) $\sigma_{k}(C_{j})=\sum_{\mathbf{x}_{i}\in C_{j}}{\lambda_{k,max}^{C_{j}}(\mathbf{x}_{i})-\lambda_{k,min}^{C_{j}}},$
where $\lambda_{k,max}^{C_{j}}(\mathbf{x}_{i})$ is the density at which $\mathbf{x}_{i}$ falls out of $C_{j}$ or $C_{j}$ separates into two clusters, and $\lambda_{k,min}^{C_{i}}$ is the minimum density at which $C_{j}$ exists. In words, a cluster’s stability is the sum of the density ranges in which points are part of the cluster. In the case of a continuous 1D density function, the clusters’ relative stabilities can be visualised as the area under the curve between the births and deaths of the clusters (see Figure 2).
@@
HDBSCAN* provides two strategies for selecting a ‘flat’ clustering: the excess of mass (eom) strategy and the leaf strategy (Campello et al., 2015). The eom strategy locally decides between selecting a cluster or its children in the condensed cluster hierarchy based on their relative stabilities. The resulting set of clusters maximises the relative cluster stability while preventing any data point from being a member of more than one selected cluster. The leaf strategy selects all leaf segments in the condensed hierarchy, typically resulting in more and smaller clusters.
Figure 2. An illustration of the relative cluster stability for a continuous 1D density function. The tree in black indicates the (condensed) cluster hierarchy or density contour tree. The coloured areas indicate each contour cluster’s relative stability. Image inspired by (McInnes and Healy, 2017).
A 1D density curve with three local maxima. The area under the curve between the local maxima and local minima is coloured, distinguishing each cluster. A cluster hierarchy tree is drawn on top, emphasising where the clusters merge.
Figure 2. An illustration of the relative cluster stability for a continuous 1D density function. The tree in black indicates the (condensed) cluster hierarchy or density contour tree. The coloured areas indicate each contour cluster’s relative stability. Image inspired by (McInnes and Healy, 2017).
3. Flare-sensitive HDBSCAN*
Conceptually, FLASC combines HDBSCAN* and the graph-based flare detection approach described in Section 1.2. An open source implementation of the algorithm is available [^1] https://github.com/vda-lab/pyflasc, which builds upon McInnes et al.’s HDBSCAN* implementation (McInnes et al., 2017).
3.1. The FLASC Algorithm
Algorithm 1 shows the high-level pseudocode of FLASC. The algorithm starts by evaluating a ‘flat’ HDBSCAN* clustering, keeping track of the space tree used in HDBSCAN* (McInnes et al., 2017; McInnes and Healy, 2017) to accelerate later steps. One noteworthy change from McInnes et al.’s implementation (McInnes et al., 2017) is that we give all points the zero label when a single cluster is allowed and selected and no cluster selection epsilon is applied, rather than only the points that joined in one of the cluster’s children in the condensed hierarchy. This enables FLASC to better analyse branching structures in data sets that contain a single cluster. Then, for each selected cluster $C_{j}$, a branch detection step is performed (potentially in parallel), which we explain in more detail below.
The first flare detection step is computing the cluster centrality $c(\mathbf{x}_{i})$ for $\mathbf{x}_{i}\in C_{j}$. We use a centrality metric that can be computed in $\mathcal{O}(N)$ based on the points’ distance to the cluster’s centroid $\bar{\mathbf{x}}_{C_{j}}$, i.e., the cluster’s weighted average (see Figure 3). HDBSCAN*’s cluster membership probabilities are used as weights. These distances are then reversed to compute the cluster centralities (see Figure 3):
(3) $c(\mathbf{x}_{i})=\max_{\mathbf{x}_{l}\in C_{j}}{\{d(\bar{\mathbf{x}}_{C_{j}},\mathbf{x}_{l})\}}-d(\bar{\mathbf{x}}_{C_{j}},\mathbf{x}_{i}).$
The second flare detection step is extracting the cluster approximation graph $\mathbf{G}_{k}^{C_{j}}$. Two types of approximation graphs are supported: the full approximation graph and the core approximation graph. Both types contain a vertex for each point in the cluster $\mathbf{x}_{i}\in C_{j}$. The approaches differ in which edges they include. The full approximation graph adds all edges with $d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{l})\leq d_{max}^{C_{j}}$, where $d_{max}^{C_{j}}$ is the largest distance in the cluster’s minimum spanning tree (MST). The resulting graph accurately describes the connectivity within the cluster at the density where the last point joins the cluster. The space tree constructed by HDBSCAN* is used to retrieve these edges efficiently. The core approximation graph adds all edges with $d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{j})\leq\max{\{\kappa{(\mathbf{x}_{i})},\kappa{(\mathbf{x}_{j})}\}}$. The resulting graph accurately describes the connectivity in the cluster’s MST. Alternatively, this graph can be considered as the cluster’s subgraph from the $k$-nearest neighbour graph over the entire data set. HDBSCAN* already extracted these edges when the core distances were computed, so this approach has a lower additional cost. Finally, the extracted edges are re-weighted with the maximum centrality of the points they connect: $\max{\{c(\mathbf{x}_{i}),c(\mathbf{x}_{l})\}},$ to form a cluster centrality graph $\mathbf{G}_{c}^{C_{j}}$ from which branches can be detected (see Figure 3).
Figure 3. A visual explanation of FLASC. (a) A centroid (red pentagon) within each cluster detected by HDBSCAN* is computed as the weighted average data point. (b) The points’ reversed distance to that centroid is used as centrality metric. (c) Single Linkage Clustering and HDBSCAN*’s hierarchy simplification is applied to the cluster’s centrality graph, resulting in a condensed branch hierarchy.
Three plots are shown in this figure. The first plot shows the data points of a Y-shaped cluster. The cluster’s weighted average data point is indicated. The second plot shows the same cluster’s data points coloured by the points’ centrality. The third plot shows three thresholded versions of the cluster centrality graph where increasingly more eccentric edges are removed, indicating how branches separate and become detectable as clusters.
Figure 3. A visual explanation of FLASC. (a) A centroid (red pentagon) within each cluster detected by HDBSCAN* is computed as the weighted average data point. (b) The points’ reversed distance to that centroid is used as centrality metric. (c) Single Linkage Clustering and HDBSCAN*’s hierarchy simplification is applied to the cluster’s centrality graph, resulting in a condensed branch hierarchy.

### v2

@@ Abstract
Clustering algorithms are often used to find subpopulations in exploratory data analysis workflows. Not only the clusters themselves, but also their shape can represent meaningful subpopulations. In this paper, we present FLASC, an algorithm that detects branches within clusters to identify such subpopulations. FLASC builds upon HDBSCAN*—a state-of-the-art density-based clustering algorithm—and detects branches in a post-processing step that describes within-cluster connectivity. Two variants of the algorithm are presented, which trade computational cost for noise robustness. We show that both variants scale similarly to HDBSCAN* in terms of computational cost and provide stable outputs using synthetic data sets, resulting in an efficient flare-sensitive clustering algorithm. In addition, we demonstrate the benefit of branch-detection on two real-world data sets.
keywords:
@@
Exploratory data analysis, density-based clustering, branch-hierarchy detection, HDBSCAN*
This work is licensed under a Creative Commons Attribution-NonCommercial-NoDerivatives 4.0 International License. This work has been submitted to PeerJ for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
Introduction
Exploratory Data Analysis (EDA)—i.e., searching for interesting patterns in data—is ubiquitous in data science and knowledge discovery workflows. Detecting which subpopulations exist in a data set is a common step in EDA. Typically, subpopulations are detected as clusters, which traditional algorithms model as Gaussian distributions (as cited in Campello et al. (2015)). More recently, density-based clustering has become more popular (e.g., Ester et al. (1996); Campello et al. (2013)). These algorithms, informally, specify clusters as regions of high density separated by regions of lower density, allowing them to capture cluster shapes, which may be non-convex and reveal relevant subpopulations. For example, a Y-shaped cluster might represent an evolving process with two distinct outcomes. Consequently, the branches in a cluster’s manifold—i.e., flares—can represent meaningful subpopulations (see also, Lum et al. (2013); Skaf and Laubenbacher (2022); Kamruzzaman et al. (2018); Reaven and Miller (1979)).
@@ Introduction
Clustering algorithms generally cannot detect this type of subgroup because there is no gap that separates flares from their cluster. From a topological perspective, clustering algorithms describe the connected components in a simplicial complex of the data (Carlsson, 2014): a set of points, edges, and triangles that describe connectivity. Flares are connected in the simplicial complex. In other words, there is a path between data points in different branches that exclusively goes through data points ‘that lie close together’. Therefore, they have a vanishing homology and cannot be detected as clusters.
Several flare-detection techniques have been proposed in topological data analysis literature. For example, Carlsson (2014) proposed functional persistence to distinguish flares from a data set’s central core. This technique quantifies data point centrality as the sum of its distances. Central observations have lower distance sums than points towards the extreme ends of the feature space. A manually controlled centrality threshold then removes the data’s core, separating branches from each other and making them detectable as clusters.
@@
Extending this approach to compute branching hierarchies requires considering both the centrality and the data point distances in what is called a bi-filtration. The centrality controls how much of the core is retained to describe how branches grow and merge. The data point distances determine whether points are connected and form a cluster. Algorithms for computing bi-filtrations are computationally expensive (Lesnick and Wright, 2022; Kerber and Rolle, 2021). Their resulting bi-graded hierarchies are also complicated to work with, as they do not have a compact representation (Carlsson, 2014) (though research into usable representations is ongoing (Botnan et al., 2022)), and existing visualisations are non-trivial (Lesnick and Wright, 2015; Scoccola and Rolle, 2023). Alternative strategies that simultaneously vary both dimensions in a single-parameter filtration exist (Chazal et al., 2009); however, they remain computationally expensive (Vandaele et al., 2021).
@@
In the present paper, we present an approach that efficiently computes branching hierarchies and detects branch-based subgroups of clusters in unfamiliar data. Inspired by Vandaele et al. (2021), we compute branching hierarchies using graph approximations of the data. This effectively replaces functional persistence’s manual centrality threshold with the question of which data points should be connected in the approximation graph. We will use HDBSCAN* (Campello et al., 2013, 2015)—a state-of-the-art density-based clustering algorithm—to answer this question. Conceptually, our approach can be thought of as creating a sequence of subgraphs which progressively include more and more central points and tracking the remaining connected components. Interestingly, a similar method has been used by Li et al. (2017) to detect actual branches in 3D models of plants.
Our main contribution is this flare detection approach, implemented as a post-processing step in McInnes et al. (2017)’s HDBSCAN* implementation111hdbscan: https://github.com/scikit-learn-contrib/hdbscan and as a stand-alone package222pyflasc: https://github.com/vda-lab/pyflasc. We propose two types of approximation graphs that naturally arise from HDBSCAN*’s design and provide a practical centrality metric for points within a cluster that is computable with a linear complexity. Combining density-based clustering and flare detection into a single algorithm provides several attractive properties:
  • The ability to detect clusters and their branches.
  • No manual thresholds, instead intuitive minimum cluster and branch sizes control the detected subgroups.
  • Low computational cost compared to multi-parameter persistence and other structure learning algorithms.
  • High branch-detection sensitivity and noise robustness by operating on HDBSCAN*-clusters, which suppress spurious noisy connectivity.
  • Branch-detection at multiple distance scales because each cluster has its own approximation graph.
We call the resulting algorithm flare-sensitive clustering (FLASC) and empirically analyse its computational cost and stability on synthetic data sets to show that the flare detection cost is relatively low. In addition, we demonstrate FLASC on two real-world data sets, illustrating its benefits for data exploration.
The remainder of this article is organised as follows: Section ‘Related Work’ provides a literature overview of related data analysis algorithms and describes the HDBSCAN* algorithm in more detail. Section ‘The FLASC Algorithm’ describes how FLASC builds on HDBSCAN* to detect branches within clusters and discusses the algorithm’s complexity and stability. Section ‘Experiments’ presents our empirical analyses that demonstrate the algorithm’s computational complexity, stability, and benefits for data exploration. Finally, Sections ‘Discussion’ and ‘Conclusion’ discuss our results and present our conclusions.
Related Work
The purpose of our work is detecting branching structures within clusters. As such, our work relates to manifold and structure learning algorithms in general. In this section, we provide an overview of related data analysis algorithms and introduce HDBSCAN*, the density-based clustering algorithm we build upon.
Structure Learning
Many types of data lie not just on a manifold, but on a smooth, one-dimensional structure. Extracting these structures can be essential in unsupervised learning applications. For example, road networks can be extracted from GPS measurements (Bonnaire et al., 2022) and cell developmental trajectories can be extracted from gene expression data (Qiu et al., 2017; Vandaele et al., 2020). Algorithms for extracting such structures are related to our work because the branch-based subgroups we are interested in can be extracted from them by partitioning the data between their intersections (Chervov et al., 2020).
Most work on extracting smooth, one-dimensional manifolds is based on Hastie and Stuetzle (1989)’s concept of principal curves: a smooth, self-consistent curve that passes through the middle of the data. Techniques estimating principal curves, trees, or graphs are often based on Expectation Maximisation (Dempster et al., 1977) and optimise the one-dimensional manifold directly (e.g., Bonnaire et al. (2022); Mao et al. (2017)). Alternative approaches are more closely related to non-linear Dimensionality Reduction (DR) algorithms that model the data’s structure as an undirected graph (e.g., Roweis and Saul (2000); Tenenbaum et al. (2000); Belkin and Niyogi (2003); van der Maaten and Hinton (2008); McInnes et al. (2020)). For example, Vandaele et al. (2020) use (manually) pruned minimum spanning trees over edges weighted by their boundary coefficient to extract a graph’s backbone. Alternatively, Ge et al. (2011) extract graph skeletons using a Reeb Graph. Reeb Graphs track the existence, merges, and splits of connected components in level sets of a continuous function defined on a manifold. Using geodesic distances to an arbitrary eccentric point as the continuous function makes the Reeb Graph capture the manifold’s skeleton. Interestingly, Mapper—an algorithm that approximates Reeb Graphs (Singh et al., 2007)—has also been used for detecting branch-based subpopulations (Kamruzzaman et al., 2018).
There are several similarities between these methods and our work. Like Vandaele et al. (2020), our approach detects tree-based branching hierarchies. Like Ge et al. (2011), our approach is topologically inspired. Where they create a Reeb Graph, we compute a Join Tree. The main difference of these methods compared to our work is their goal. We aim to identify relevant branch-based subpopulations. Ge et al. (2011), Vandaele et al. (2020), and the expectation maximisation-based algorithms explicitly model the data’s structure which necessitates a larger computational cost.
HDBSCAN*
(a) 2D data
(b) Density profile
(c) HDBSCAN* clusters
(d) Cluster hierarchy
Figure 1: Density-based clustering concepts behind HDBSCAN*. (a) A 2D example point cloud with varying density adapted from McInnes et al. (2022)’s online tutorial. (b) Density contours in a height map illustrate the data’s density profile. Peaks in this density profile correspond to density contour clusters. (c) Clusters extracted from the density profile by HDBSCAN* indicated in colour. (d) The density contour tree describes how density contour clusters merge when considering lower density thresholds.
HDBSCAN* is a density-based clustering algorithm that provides state-of-the-art clustering performance (Campello et al., 2013, 2015). Informally, density-based clustering specifies clusters as regions of high density separated by regions of lower density. This formulation does not limit clusters to convex shapes and provides a natural way to separate noise points from clusters. The algorithm is well suited for exploring unfamiliar data because—unlike older popular clustering algorithms—HDBSCAN* does not require the number of clusters or the distance between clusters to be specified in advance.
Several studies have implemented and adapted the HDBSCAN* algorithm: McInnes and Healy (2017) improved the algorithm’s computational performance by using space trees for finding the data points’ nearest neighbours and provide a popular and efficient Python implementation (McInnes et al., 2017). Stewart and Al-Khassaweneh (2022) created a Java implementation with a novel prediction technique for unseen data points. Jackson et al. (2018) presented an approximate HDBSCAN* algorithm that uses NN-descent (Dong et al., 2011) for finding the nearest neighbours providing fast distributed performance. Malzer and Baum (2020) introduced a cluster selection distance threshold that effectively creates a hybrid between DBSCAN’s (Ester et al., 1996) and HDBSCAN*’s cluster selection, improving the algorithm’s performance on data sets with small clusters and a large density variability. Neto et al. (2021) showed how Relative Neighbourhood Graphs (RNGs) (Toussaint, 1980) can be used to efficiently compute HDBSCAN* cluster hierarchies for multiple min cluster size values. Their follow-up work presented MustaCHE, a visualisation tool for the resulting meta-cluster hierarchy (Neto et al., 2018). To our knowledge, no previous study has adapted HDBSCAN* for detecting flares.
@@
Because our work builds on HDBSCAN*, it is relevant to explain how the algorithm works in more detail. The remainder of this section describes HDBSCAN* following Campello et al. (2015)’s explanation. We refer the reader to McInnes and Healy (2017) for a more formal, statistically motivated description of the algorithm.
The HDBSCAN* Algorithm
(a) Cluster centroid
(b) Eccentricity profile
(c) FLASC branches
(d) Branch hierarchy
Figure 2: Density-based clustering concepts behind FLASC. (a) A within-cluster eccentricity $e(\mathbf{x}_{i})$ is defined for each point $\mathbf{x}_{i}$ in cluster $C_{j}$ based on distances to the cluster’s membership weighted average shown by the pentagon mark. (b) The cluster’s eccentricity profile visualised as contours on a height map. Peaks in the profile correspond to branches in the cluster. (c) Branches extracted from the cluster by FLASC indicated in colour. The cluster’s centre is given its own label. (d) The eccentricity contour tree describes how branches merge when considering lower eccentricity thresholds.
HDBSCAN* is based on density-based clustering concepts pioneered by Wishart (1969) and formalised by Hartigan (1975). We demonstrate these ideas using a 2D point cloud adapted from McInnes et al. (2022)’s online tutorial shown in Fig. 1(a). In general, let $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ be a data set consisting of $N$ feature vectors $\mathbf{x}_{(\cdot)}$ and a distance metric $d(\mathbf{x}_{i},\mathbf{x}_{j})$. Then, a point’s core distance $\kappa(\mathbf{x}_{i})$ is the distance to its $k$-nearest neighbour and HDBSCAN* estimates its density as $\lambda_{k}(\mathbf{x}_{i})=1/\kappa(\mathbf{x}_{i})$ (Campello et al., 2015). Figure 1(b) illustrates the example’s density profile as contours in a height map. Density contour clusters intuitively correspond to peaks in the density profile, for example, the clusters indicated in colour in Fig. 1(c). More formally, the density contour clusters at some threshold $\lambda_{t}$ are a collection of maximal, connected subsets in a level set $\{\mathbf{x}\;|\;\lambda(\mathbf{x})\geq\lambda_{t}\}$ (Hartigan, 1975). In other words, density contour clusters are the connected components of points with a density larger than some threshold. Density contour trees capture the hierarchy in which density contour clusters merge as the density threshold decreases. From a topological perspective, density contour trees are a join tree of the data’s density profile.
Data sets generally do not have an inherent notion of connectivity between their data points. Such a notion is needed to determine whether two points are part of the same density contour cluster at a threshold $\lambda_{t}$. HDBSCAN* solves this problem by considering points to be connected if the distance between them is smaller than or equal to $1/\lambda_{t}$. This solution is possible because density is defined in terms of distance. HDBSCAN* uses a mutual reachability distance between points for this purpose, which is defined as (Campello et al., 2015):
@@
$d_{mreach}(\mathbf{x}_{i},\mathbf{x}_{j})=\begin{cases}\max{\left\{\kappa(\mathbf{x}_{i}),\kappa(\mathbf{x}_{j}),d(\mathbf{x}_{i},\mathbf{x}_{j})\right\}}&\text{if }\mathbf{x}_{i}\neq\mathbf{x}_{j}\text{,}\\ 0&\text{otherwise,}\end{cases}$ (1)
where the value of $k$ as used in $\kappa$ is specified manually and acts as a smoothing factor for the density estimation.
@@
Using the mutual reachability distance to provide connectivity, we can now recover the density contour tree. The edges that change connectivity between density contour clusters are exactly those edges in the data’s Minimum Spanning Tree (MST)  (see, f.i., Doraiswamy et al. (2021)). HDBSCAN* uses a MST to efficiently compute a single linkage clustering hierarchy (Sibson, 1973). The resulting dendrogram is simplified using a manually specified minimum cluster size $m_{c}$ to recover a condensed cluster hierarchy that resembles the data’s density profile as shown in Fig. 1(d). From the root down, only the sides of a split containing more than $m_{c}$ points are considered to represent clusters. Sides with fewer points are interpreted as “falling out of the parent cluster” (McInnes and Healy, 2017) or the cluster disappearing completely.
@@
HDBSCAN* provides two strategies for selecting clusters from the condensed hierarchy: the excess of mass (EOM) strategy and the leaf strategy (Campello et al., 2015). The leaf strategy selects all leaf segments in the condensed hierarchy, typically resulting in multiple small clusters. The EOM strategy maximises relative cluster stability while preventing any data point from being a member of more than one selected cluster. A cluster $C_{j}$’s relative stability $\sigma_{k}(C_{j})$ is defined as (Campello et al., 2015):
$\sigma_{k}(C_{j})=\sum_{\mathbf{x}_{i}\in C_{j}}{\lambda_{k,max}^{C_{j}}(\mathbf{x}_{i})-\lambda_{k,min}^{C_{j}}},$ (2)
where $\lambda_{k,max}^{C_{j}}(\mathbf{x}_{i})$ is the density at which $\mathbf{x}_{i}$ falls out of $C_{j}$ or $C_{j}$ separates into two clusters, and $\lambda_{k,min}^{C_{i}}$ is the minimum density at which $C_{j}$ exists. In words, the stability of a cluster is the sum of density ranges in which points are part of the cluster, which corresponds to the area of the cluster’s icicle in Fig. 1(d). HDBSCAN*’s cluster selection epsilon parameter can be used to specify a minimum persistence for EOM clusters (Malzer and Baum, 2020).
Flare-sensitive HDBSCAN*